OpenStudy (anonymous):

3log2^5 = x and log2^5^3 = x are the same equation. TRUE OR FALSE

OpenStudy (anonymous):

true

OpenStudy (anonymous):

what happens when you raise an exponent to an exponent?

OpenStudy (anonymous):

this is true because using property of log (exponents ) gets down so \[\Large \implies \log _{a}^{b^n}=nlog _{a}^{b}\] so \[\Large \implies \log _{2}^{5^3}=3\log _{2}^{5}\]

OpenStudy (anonymous):

log2^5 + log2^6 = log2^11 true or false

OpenStudy (anonymous):

http://www.purplemath.com/modules/logrules.htm

OpenStudy (anonymous):

thats a good side

OpenStudy (anonymous):

so it true

OpenStudy (anonymous):

yes it is. look at the above !